How high from the floor is the spot on the wall?

A light ray strikes a plane mirror lying flat on a horizontal floor, making an angle of incidence = 61.8°. After reflecting from the mirror, the ray makes a spot on a vertical wall. The wall is a horizontal distance d=1.08 m from the point of reflection, as shown. How high from the floor is the spot on the wall? (Neglect the thickness of the mirror.)
Answer #1

BY the law of reflection

Angle of reflection = angle of incidence

= 64.2o

angle that the lower makes with the horizontal = 90 - 64.2 = 25.8o

tan 25.8 = h/d

h = d tan 25.8

= 1.08*tan 25.8

= 0.522 m


answered by: ANURANJAN SARSAM